Q. A company with $65500 in current assets and $37000 in current liabilities pays a $1900 current liability. As a result of this transaction, the current ratio and working capital will

increase and remain the same, respectively.

both decrease.

both increase.

remain the same and decrease, respectively.
Answer #1

Solution:

Existing current ratio = 65500/ 37000 = 1.77

Existing working capital = 65500 - 37000 = $28,500

New current ratio = (65500-1900) / (37000-1900)= 1.81

New working capital = (65500-1900) - (37000-1900) = $28,500

Therefore, current ratio will increase and working capital will remain the same.

Hence first option is correct.